Child's Play Company makes a plastic rattle for toddlers. The rattle is generally marketed through exclusive retailers located in upscale shopping malls. The company used 23 retailers. In late 2016, Diana Suares, the president of the company, was considering alternative marketing plans presented to her by Bill Dings, the marketing manager. Based on sales from January through October 2016, Diana expected that 2016sales would amount to 300,000 units at $8.00 per unit.

Diana Suares also had with her some cost data for 2016 supplied by the CFO, Don Capp. Don expects that these costs are reliable estimates for a volume up to 400,000 units. Beyond 400,000 units, the company would have to rent additional machines (with a capacity of 100,000 units each) at an annual cost of $50,000 per machine (in 2016). The data are presented in Exhibit A.

Total fixed manufacturing and selling and administrative costs are each expected to increase in 2017 by 10 percent because of inflation. Variable costs per unit and the selling price would stay the same as in 2016.

Manufacturing costs for rattles (based on production volume of 300,000 units).

Direct material $0.80 per unit *

Direct labor $10 per hour * (Each worker can make 20 units in an hour)

Packaging $0.75 per unit *

Power, supplies, indirect labor, and other variable production costs $1.20 per unit *

Supervisory salaries, equipment rental, and miscellaneous production costs $1.80 per unit *

These costs vary in total with production volume Selling and administrative costs (based on sales volume of 300,000 units).

Commissions to sales staff 10% of price to retailer **

Shipping $0.50 per unit **

Advertising and promotion $0.60 per unit

Administrative staff salaries, depreciationon office equipment, etc $0.90 per unit **

These costs vary in total with sales volume.

What is the contribution margin per unit?

What is the contribution margin if 1000 units are sold?

What is the contribution margin if100,000 units are sold?

What is the profit at the break even volume?

What is the contribution margin at the break even volume?

What is the break even volume?

What is the profit at 1000 units above break even?

What is the profit at 2000 units below the break even?

What is the profit at 100,000 units above break even?

What is the margin of safety at 1000 units above break even?

What is the sales volume required to give $100,000 profit before tax?

What is the sales volume required to give $180,000 after tax, assuming a tax rate of 40%?
Answer #1

In case more than 4 questions are asked only the first four are answered:-

1. What is the contribution margin per unit?

Contribution margin = contribution per unit /selling price per unit.

Contribution per unit = selling price per unit - variable cost per unit

Selling price per unit = $8.00

Variable cost per unit = Direct material $0.80 per unit + Direct labor $0.50 per unit (Each worker can make 20 units in an hour, rate per hour is $ 10 thus rate per unit =10/20 ) + Packaging $0.75 per unit + Power, supplies, indirect labor, and other variable production costs $1.20 per unit.

Variable cost per unit = $3.25

Contribution = 8.00 - 3.25 = $ 4.75

Contribution margin = 4.75 / 8 = 60%.

2. What is the contribution margin if 1000 units are sold?

3. What is the contribution margin if 100,000 units are sold?

Answer for all 3 question is the same, since the variable cost per unit would not change for any number of units produced, the contribution per unit would also not change, similarly the contribution margin would also not change.

4. What is the profit at the break even volume?

Break even point is where the total contribution is equal to the total fixed cost, i.e the total sales = total cost and there is zero profit. Therefore there would be no profit at break even volume.